Active Incident

Updated 13 hours ago

QLD Packet LossOperational

Incident Status

Operational

Components

Peering, Intercapital Links, Fibre Services, Ethernet Services, DSL/NBN Services



May 10, 2021 11:45AM AEST
[Identified] Data Express NOC has identified a problem within our third-party carrier network causing some packet loss. The issue is affecting some services in Queensland. Data Express is working with the carrier to resolve this issue. DX NOC 1300 801110

May 10, 2021 3:13PM AEST
[Identified] - Issue is currently under investigation with the upstream providers.

May 10, 2021 4:12PM AEST
[Monitoring] The upstream provider has resolved the issue. DX will provide a PIR once all details have been received.

Incident Status

Operational

Components

Core Systems



March 5, 2021 6:12AM AEDT
[Monitoring] Dear valued customer, At approximately 2:00am this morning (05/03/2021 AEDT), DX Engineers were notified of increased traffic to the CloudExpress platform. We are currently investigating the origin of the traffic and have taken steps to minimise the impact on the platform. Impacted Services: DX CloudExpress Hosting Services. DX NOC 1300 801110

March 5, 2021 8:51AM AEDT
[Monitoring] DDoS attack confirmed against the CloudExpress platform. Attack was aimed at one of our CloudExpress customers. DDoS mitigation immediately kicked in and isolated the traffic within the EDGE (Border Router network) in NSW. Due to the multi-homed nature of the DX network, traffic from some networks (Isolated to one IP transit provider) to the DX network may have been impacted whilst traversing to the CloudExpress platform, this only affected a small portion of inbound traffic to DX NSW Core. Updates to come. DX NOC 1300 801110

March 5, 2021 3:01PM AEDT
[Monitoring] We are still seeing a large amount of traffic hitting this platform. DX Engineers are re-routing traffic to larger capacity DDoS scrubbers upstream.

March 5, 2021 5:40PM AEDT
[Monitoring] DX Engineers have worked with the clients affected and can confirm that the platform is now stable. DX will closely monitor the situation over the next 72 hours for any re-occurring attacks. DX NOC 1300 801110
NSW - Services DownPartial Service Disruption

Incident Status

Partial Service Disruption

Components

Fibre Services, Ethernet Services, DSL/NBN Services



February 25, 2021 10:30AM AEDT
[Investigating] Data Express NOC has identified an outage affecting some services in NSW. Fibre services with DualWAN should failover to the secondary service. The issue is still under investigation, and DX NOC is working closely with the carriers and 3rd party providers to resolve the issue. There is no ETR at this stage. Further updates will be provided within an hour. DX NOC - 1300801110

February 25, 2021 10:50AM AEDT
[Monitoring] The third party carrier has identified and resolved the issue. All services are being restored. Further updates to follow. Please contact Data Express for more information.

Incident Status

Operational

Components

Peering, Intercapital Links, Fibre Services, Ethernet Services



February 17, 2021 9:26PM AEDT
[Investigating] Data Express NOC has identified an outage affecting some services in Victoria. Data Express is working to further isolate the issue. DX NOC 1300 801110

February 17, 2021 9:37PM AEDT
[Monitoring] All services stable, Engineers are investigating the cause.
NSW - Fibre Services DownPartial Service Disruption

Incident Status

Partial Service Disruption

Components

Fibre Services



January 13, 2021 3:47PM AEDT
[Investigating] Data Express NOC has identified an outage affecting some services in NSW. Fibre services with 4G backup should failover to the backup link. The issue is still under investigation, and DX NOC is working closely with the carriers and 3rd party providers to resolve the issue. There is no ETR at this stage. Further updates will be provided within an hour. DX NOC - 1300801110

January 13, 2021 9:23PM AEDT
[Monitoring] All services being restored now. More updates to follow once the 3rd party restores the issue from their side. If you are still experiencing an issue, please contact DX NOC on 1300801110.

January 14, 2021 12:31PM AEDT
[Monitoring] Our upstream vendor who is currently working on the issue we encountered yesterday (primary fibre port failure), reconfigured and issued a command which caused a large storm within our routing cluster at Sydney SY3. This caused an outage for all Sydney services within the DX Sydney POP this morning (10:54 am - 11:53 am.) DX Engineers rectified the issue and have put in place safeguards to stop this issue occurring again. More information to come once we have it and a PIR will be released once we receive it.

January 14, 2021 1:41PM AEDT
[Monitoring] More services restored back to Fibre. There are a few sites in specific route zones that have not failed back to Fibre from LTE, we are working with our vendor to identify and fix the issue asap. Updates to come.

IP Transit




Operational

Peering




Partial Service Disruption

Intercapital Links




Partial Service Disruption

Fibre Services




Operational

Ethernet Services




Partial Service Disruption

DSL/NBN Services




Partial Service Disruption

Voice Services




Operational

Core Systems




Operational

External Services

MS Office365

SoftLayer